214.2 gm of aquous solution of sugar which contains 34.2 gm sugar calculate the mass of solution

Grade:12

1 Answers

Arun
5 years ago
(i) Weight of sugar syrup = 214.2 g
Weight of sugar in the syrup = 34.2 g
∴ Weight of water in the syrup = 214.2 – 34.2 = 180.0 g
Mol. wt. of sugar, C12H22O11 = 342
∴Molal concentration = 34.2 *1000/342 *180 = 0.56
(ii) Mol. wt. water, H2O = 18
∴ Mole fraction of sugar = 34.2/342 /180/18+34.2/342
= 0.1/10+0.1 = 0.1/10.1 = 0.0099